# SPSysML: A meta-model for quantitative evaluation of Simulation-Physical Systems

### v3

@@ 6 Ilustrative example - goal validation
The robot in the INCARE project coexists with humans; thus, we utilise WorldSync <<WorldMirrGpAgents>> to manage humans in the Simulated World. Detailed specification and realisation of WorldSync <<WorldMirrGpAgents>> is available in [49].
Step 4 and 5
In the final design iteration, the system is decomposed to 5 embodiment-specific Agents, one for each system part: TIAGo <<SimAgent>> mirroring TIAGo <<PhyAgent>>, FallDetector <<SimAgent>> mirroring FallDetector <<PhyAgent>> and SmartHome <<PhyAgent>>. Additionally, there are 3 Hybrid Agents, one for each computational function of the system: ComplexTaskExecution <<HybAgent>>, Talker <<HybAgent>>, FakeAudio <<HybAgent>>. TIAGo <<SimAgent>>, FakeAudio <<HybAgent>> and Talker <<HybAgent>> compose Robot <<MirrSimGpAgents>>, and TIAGo <<PhyAgent>> and Talker <<HybAgent>> compose Robot <<MirrPhyGpAgents>>.
Step 6
@@
To provide an example, we describe the final decomposition of two TIAGo <<SimAgent>> realisations (O3deTIAGo (Fig. 15) being the robot simulator implemented in the O3DE simulator and GazeboTIAGo (Fig. 16) implemented in Gazebo) and TIAGo <<PhyAgent>> (Fig. 17). Each robot hardware component is specified as either <<PhyRealRec>>, <<PhyRealEff>>, <<SimRealRec>>, or <<SimRealEff>>. O3deTIAGo: TIAGo <<SimAgent>> integrates the system with O3DE simulation environment; thus, <<SimRealRec>>s and <<SimRealEff>>s are core O3DE components interacting with the simulated world, called gems. GazeboTIAGo: TIAGo <<SimAgent>> integrates the system with Gazebo simulation environment; thus, <<SimRealRec>>s and <<SimRealEff>>s expose gazebo::SensorPlugin and gazebo::ModelPlugin interfaces accordingly. As the physical robot, we use PAL Robotics’ TIAGo; thus, <<PhyRealRec>>s, <<PhyRealEff>>s interfaces are adequate Linux Kernel drivers managing communication with the devices. In GazeboTIAGo: TIAGo <<SimAgent>> case, the Simulated Drivers connect to the gazebo::SensorPlugin and gazebo::ModelPlugin interfaces and expose the robot state information and typical ROS topics/services (e.g. JointStateInterface and EffortJointInterface for MobileBaseController <<SimVirtEff>> [^5] This ¡¡VirtEff¿¿ is based on gazebo_ros_control package: https://classic.gazebosim.org/tutorials?tut=ros_control and /scan ROS topic for lidar <<SimVirtRec>>). If Simulated World is the Gazebo environment, <<SimVirtEff>>s and <<SimVirtRec>>s are usually implemented as Gazebo Plugins. <<PhyDriverGpSubsys>> connects to Linux Kernel drivers and, as a whole Group, exposes to <<ContSubsys>> identical interfaces as <<SimDriverGpSubsys>>. The diagrams of the TIAGo robot agents (simulated and physical) show their common RobotIf <<SimPhyContSubsys>>; however, as the physical robot and Gazebo simulator run ROS1: RobotIf <<SimPhyContSubsys>>, and O3DE simulator runs ROS2: RobotIf <<SimPhyContSubsys>>.
@@
Step 7
[…]
Step 8
Based on the <<OptionalReq>>s in the requirements we define 12 <<SysSetupGpAgents>>s for (SmartHome <<PhyAgent>> existing or not, FallDetector <<PhyAgent>> or FallDetector <<SimAgent>> or no FallDetecor, Robot <<PhyAgent>> or Robot <<SimAgent>>). We show Simulated <<SysSetupGpAgents>> (Fig. 18) and Physical <<SysSetupGpAgents>> (Fig. 19) Internal Block Diagrams (IBDs) to exemplify <<SysSetupGpAgents>> specification. Simulated <<SysSetupGpAgents>> presents two realisation of TIAGo <<SimAgent>>GazeboTIAGo and O3deTIAGo.
@@
Step 9 and 10
In INCARE the Robot <<PhyAgent>> and Robot <<SimAgent>> are realised with TIAGo robot and its simulations. PAL Robotics mostly implemented these. We mapped the robot hardware and software to our design (as shown in Fig. 15, 16 and 17). RosControl <<SimDriverGpSubsys>> is a software package [^6] http://wiki.ros.org/ros_control implementing various common ROS controllers, and SimTIAGoDrives <<SimHardGpSubsys>> is a package [^7] http://wiki.ros.org/pal_hardware_gazebo implementing Gazebo plugins controlling TIAGo drives. Additionally, we boosted the robot’s voice communication ability using the Google Dialogflow service. We equipped TIAGo <<PhyAgent>> with additional USB Microphones [55] to improve its audio perception. The other <<Agent>>s were implemented and tested following the simulation implementation factor. In the result, we got Simulated <<SysSetupGpAgents>> implemented and tested before Physical <<SysSetupGpAgents>>, what improved robot system development safety. The simulated and physical worlds’ synchronisation feature (mirroring humans in the Simulated World) was implemented with HuBeRo framework as WorldSync <<WorldMirrGpAgents>> and tested in a crowded hospital environment (Fig. 20).
@@
The INCARE system was deployed in an end-user home, and the videos present its performance in the example tasks— transportation (Fig. 2(a)) [^8] https://vimeo.com/670252925 and fall assistance (Fig. 2(b)) [^9] https://vimeo.com/670246589.
7 Summary
Numerous Cyber-Physical Systems include simulation parts either as Digital Twins (DT), demonstrators, or mockups utilised during their development. We postulate the Simulation-Physical System (SPSys) concept to describe this kind of system. SPSys includes physical, simulated and hybrid parts. They cooperate to fulfil the system’s aim; however, some are used as DTs to boost the system’s reliability and analysis possibility. Some parts of the system can be used in the development process only as mockups and prototype simulators to increase the simulation-based testing coverage of the system. This, in turn, improves prototype safety and the resulting system robustness.
SPSys application is wide; in particular, it can work in a dynamic environment so that it can observe exogenous actions of the environment’s habitants. Such a situation is problematic because DT must perfectly mirror its PT. To answer the above needs, we propose a Domain-Specific Language named Simulation-Physical System Modeling Language (SPSysML) that models the SPSys artefact and demonstrates SPSys taxonomy and the relationships between the types of SPSys parts.
One of the crucial aspects of reliable software development is integrity and reusability. Therefore, we propose an SPSys Development Procedure (SPSysDP) that, using design evaluation factors, supports quantitative analysis and optimisation targeted to software re-use maximisation between DT and PT and in different system setups. We analyze the evaluation factors and show features of the systems that score edge case values of the factors.
@@ 7 Summary
Finally, we verify SPSysML and SPSysDP in complex robot system development. We demonstrate step-by-step SPSysDP execution. We point out significant system structure changes resulting from the design evaluation and the proposed quantitative factor-based guidelines applications. An example is FallDetector <<SimAgent>> and FallDetector <<PhyAgent>> integration improvement by common part isolation as <<SimPhyContSubsys>>. Moreover, the verification shows that 33% of the system’s Physical Agents do not have DTs, and thanks to the design modification, the whole hardware (physical and simulated) is controlled with <<SimPhyContSubsys>> (DGF=1). This has two main advantages. First, all Digital and Physical Twins pairs share a common control subsystem. Second, if there is a <<PhyAgent>> or <<SimAgent>> without a twin, it can be easily integrated with its Digital or Physical Twin in the future.
@@
The conducted validation shows that the requirement profile for SPSys enables critical analysis and may result in requirement decomposition and explication. An example is the decomposition of the Communication with humans <<SimPhyFunReq>> to TTS and STT <<SimPhyFunReq>> and Dialoge management <<CompFunReq>> in the first and second steps of SPSysDP. Thanks to the Dialoge management <<CompFunReq>> separation, the proposed requirement-based procedure for the system decomposition resulted in the implementation of Talker <<HybAgent>> managing the requirement. As a consequence of the above, the requirement is managed by a component usable in different setups in cooperation with the simulated or physical embodiment. Otherwise, the requirement would be managed separately in simulation and physical embodiments of the robot. The presented validation confirms the features distinguishing our method from the others listed in Table I:
  • Integrity evaluation with quantitative factors,
  • SysML-based structure model and system development procedure,
  • System design optimisation considering simulation and physical embodiments of the system,
@@
  • Multiple system execution and testing setups specification,
  • Forcing simulation-based testing prior physical embodiment development using simulation implementation factor for sequencing <<SysSetupGpAgents>> implementation and testing,
@@
  • Specification of Simulation-Physical System including multiple physical and simulated parts and multiple Digital Twins.
  • <<WorldMirrGpAgents>> artefact definition and including its design in SPSysDP to reflect the dynamic environment in simulation, enabling Digital Twins to observe exogenous actions.
@@
  • SPSysDP is a roadmap for SPSys’s development. Various methodologies can be defined by setting appropriate logic predicates in decision nodes. Therefore, SPSysDP does not specify whether the system is developed top-down, bottom-up or middle-out. The final development approach depends on the decision nodes specification, and it is a challenge for a specific system. However, the literature does not specify either a universal terminal predicate for systems’ requirement definition or system designing.
  • The proposed, systematic requirement-based system structuring method is a good starting point for system design and guides the designer to associate mirroring agents and provides a straightforward association between the system components and its requirements. However, it may not be optimal for any system; therefore, we suggest further iterating the design stage with other system design methods, like those based on the Design Structure Matrix [56].
  • The requirement model defines only types utilised by the proposed requirement-based structuring method; however, other requirement types may be necessary for a specific Simulation-Physical system. In such a case the project team should extend the requirement model.
In the future, we plan to automate testing SPSysML structure and code generation based on it with known, universal tools like Matlab and Enterprise Architect. Furthermore, this work revealed the need for a CPS structure optimisation indicators taxonomy. SPSysML and future work lead to optimal structure development automation for complex robot systems.

### v5

@@ 6.1 Example system design and analysis
The robot in the INCARE project coexists with humans being exogenous agents from its perspective; thus, we utilise WorldSync <<WorldMirrGpAgents>> to manage humans in the Simulated World. Detailed specification and realisation of WorldSync <<WorldMirrGpAgents>> is available in [56].
System composition
System structuring resulted in the system being composed of 5 embodiment-specific Agents, one for each system part:
  • TIAGo <<SimAgent>>, TIAGo <<PhyAgent>>,
  • FallDetector <<SimAgent>>, FallDetector <<PhyAgent>>,
  • SmartHome <<PhyAgent>>.
Additionally, there are 3 Hybrid Agents, one for each computational function of the system:
  • ComplexTaskExecution <<HybAgent>>,
  • Talker <<HybAgent>>,
  • FakeAudio <<HybAgent>>.
There are two DT-PT pairs:
  • one resulting from the requirement Robot <<SimPhyPartReq>>:
    • Robot <<MirrSimGpAgents>> being a DT composed of TIAGo <<SimAgent>>, FakeAudio <<HybAgent>> and Talker <<HybAgent>>,
    • Robot <<MirrPhyGpAgents>> being PT composed of TIAGo <<PhyAgent>> and Talker <<HybAgent>>.
  • one resulting from the requirement Fall Detector <<OptionalReq,SimPhyPartReq>>, FallDetector <<SimAgent>> as DT and FallDetector <<PhyAgent>> as PT.
Agent decomposition
@@
To provide an example, we describe the final decomposition of two TIAGo <<SimAgent>> realisations (O3deTIAGo (Fig. 11) being the robot simulator implemented in the O3DE simulator and GazeboTIAGo (Fig. 12) implemented in Gazebo) and TIAGo <<PhyAgent>> (Fig. 13) which mirrors the previous ones. Each robot hardware component is specified as either <<PhyRealRec>>, <<PhyRealEff>>, <<SimRealRec>>, or <<SimRealEff>>. O3deTIAGo: TIAGo <<SimAgent>> integrates the system with O3DE simulation environment; thus, <<SimRealRec>>s and <<SimRealEff>>s are core O3DE components interacting with the simulated world, called gems. GazeboTIAGo: TIAGo <<SimAgent>> integrates the system with Gazebo simulation environment; thus, <<SimRealRec>>s and <<SimRealEff>>s expose gazebo::SensorPlugin and gazebo::ModelPlugin interfaces accordingly. As the physical robot, we use PAL Robotics’ TIAGo; thus, <<PhyRealRec>>s, <<PhyRealEff>>s interfaces are adequate Linux Kernel drivers managing communication with the devices. In GazeboTIAGo: TIAGo <<SimAgent>> case, the Simulated Drivers connect to the gazebo::SensorPlugin and gazebo::ModelPlugin interfaces and expose the robot state information and typical ROS topics/services (e.g. JointStateInterface and EffortJointInterface for MobileBaseController <<SimVirtEff>> [^5] This ¡¡VirtEff¿¿ is based on gazebo_ros_control package: https://classic.gazebosim.org/tutorials?tut=ros_control and /scan ROS topic for lidar <<SimVirtRec>>). If Simulated World is the Gazebo environment, <<SimVirtEff>>s and <<SimVirtRec>>s are usually implemented as Gazebo Plugins. <<PhyDriverGpSubsys>> connects to Linux Kernel drivers and, as a whole Group, exposes to <<ContSubsys>> identical interfaces as <<SimDriverGpSubsys>>. The diagrams of the TIAGo robot agents (simulated and physical) show their common RobotIf <<SimPhyContSubsys>>; however, the physical robot and Gazebo simulator run ROS1: RobotIf <<SimPhyContSubsys>>, and O3DE simulator runs ROS2: RobotIf <<SimPhyContSubsys>>.
@@
Structure evaluation
The final structure evaluation resulted with: IIF=1, MIFRobot=1, MIFFallDetector=1, DGF=1, and DTC=0.67. The result means the structure consists of no <<SimContSubsys>> or <<PhyContSubsys>>, and one <<PhyAgent>> is not mirrored by a DT— SmartHome <<SimAgent>>. The lack of DT for SmartHome <<SimAgent>> results from the requirements (Fig. 10(a)), where Smart Home is not <<SimPhyPartReq>>); thus, as the requirement is not modified, this case is the designer’s informed decision. DGF=1 means the considered SmartHome <<SimAgent>> has <<SimPhyContSubsys>>; therefore, its control subsystem can be shared with a DT, if one will be required in the future. One of the previous design iterations resulted with: IIF=$\frac{5}{7}=0.71$, MIFRobot=1, MIFFallDetector=0, DGF=$\frac{20}{22}=0.91$, DTC=0.67. In this iteration the FallDetector <<SimAgent>> and FallDetector <<PhyAgent>> use embodiment specific control layer (<<SimContSubsys>> and <<PhyContSubsys>>), because FallDetector <<SimAgent>> consists <<SimContSubsys>>, which can not be shared with its DT. To increase MIFFallDetector and DGF the common part of FallDetector <<SimContSubsys>> and FallDetector <<PhyContSubsys>> was extracted. The common part constitutes FallDetector <<SimPhyContSubsys>> in the final design. The FallDetector <<SimPhyContSubsys>> becomes a universal interface between ComplexTaskExecution <<HybAgent>> and the driver layer of FallDetector <<PhyAgent>>. It forced decomposition of FallDetector <<SimAgent>> to FallDetector <<SimPhyContSubsys>>, FallDetector <<SimVirtRec>> and FallDetector <<SimRealRec>>. The latter two simulate Fall Detector sensing. Thanks to the design evaluation, FallDetector <<SimPhyContSubsys>> is used in both embodiments in the final structure, and this part of FallDetector will be tested in simulation because it is common in simulated and physical embodiments of the system.
Configurability definition
Based on the <<OptionalReq>>s in the requirements we define 6 <<SysSetupGpAgents>>s composing of:
  • FallDetector <<PhyAgent>> or FallDetector <<SimAgent>> or no FallDetecor, and
  • Robot <<MirrPhyGpAgents>> or Robot <<MirrSimGpAgents>>.
We show Simulated <<SysSetupGpAgents>> (Fig. 14) and Physical <<SysSetupGpAgents>> (Fig. 15) Internal Block Diagrams (IBDs) to exemplify <<SysSetupGpAgents>> specification. Simulated <<SysSetupGpAgents>> presents two realisations of TIAGo <<SimAgent>>GazeboTIAGo and O3deTIAGo.
@@
Implementation
In INCARE the Robot <<PhyAgent>> and Robot <<SimAgent>> are realised with TIAGo robot and its simulations. PAL Robotics mostly implemented these. We mapped the robot hardware and software to our design (as shown in Fig. 11, 12 and 13). RosControl <<SimDriverGpSubsys>> is a software package [^6] http://wiki.ros.org/ros_control implementing various common ROS controllers, and SimTIAGoDrives <<SimHardGpSubsys>> is a package [^7] http://wiki.ros.org/pal_hardware_gazebo implementing Gazebo plugins controlling TIAGo drives. Additionally, we improved the robot’s voice communication by integrating a Large Language Model for user request interpretation [38]. We equipped TIAGo <<PhyAgent>> with additional USB Microphones [57] to improve its audio perception. Complex Task Execution <<HybAgent>> is realised with TaskER framework [55, 36]. The simulated and physical worlds’ synchronisation feature (mirroring humans in the Simulated World) was implemented with HuBeRo framework as WorldSync <<WorldMirrGpAgents>> and tested in a crowded hospital environment (Fig. 16).
@@
The INCARE system was deployed and validated in an end-user home, and the videos present its performance in the example tasks— transportation (Fig. 2(a)) [^8] https://vimeo.com/670252925 and fall assistance (Fig. 2(b)) [^9] https://vimeo.com/670246589.
6.2 SPSysML assessment
Based on the open-source SPSysML documentation, a third-party developer designed a system for quantitative evaluation of autonomous vacuum cleaner navigation algorithm [58]. This design activity scored the best possible evaluation factor values. The designer is sure the requirement-based system structuring helps in comprehensive analysis and would use it again as it simplifies CPS design. He commented on the SPSysML: SPSysML promotes wider simulation-based V&V of the system’s change, and evaluation factors help validate and point out the system’s weak parts. However, I am unsure how to implement hybrid agents in ROS, e.g., the same communication topics and structure. Early decomposition is the hardest part of the SPSysML application.
To evaluate our methodology wider, we asked systems engineering and robotics practitioners to assess SPSysML from various points of view. The questions regard individual experience Fig. 17, SPSysML utility Fig. 18, factor improvement guidance Fig. 19, drawbacks, difficulties and our methodology coupling possibilities with standard tools like Zachman Framework [59] and V-model [60]. Finally, the responders assessed our methodology against the qualities of great models Fig. 20. The key questionnaire form and all results are published [^10] https://github.com/RCPRG-ros-pkg/spsysml. The example responses include drawbacks:
  • Evaluation factors might be abused as the importance of components is not taken into account
  • The proposed design evaluation factors are not clearly understandable without a good grasp of SPSysML.
  • With a large number of requirements that have a complex structure, the diagrams may be unreadable. The solution would be a requirements association table.
and advantages:
  • SPSysML certainly gives you the opportunity to evaluate the system architecture and make conscious design decisions.
  • At early stage of system development SPSysML can show some issues regarding system design
  • Joining simulated and physical models provides numerous benefits. SPSysML allows to clearly depict this in system architecture.
  • SPSysML is a systematic framework for joining simulation and physical systems. This process can be controlled with SPSysML.
All responders would apply the requirement-based composition, and eight of nine would apply SPSysML for system specification and design evaluation.
Figure 17: Tasks in which the responders utilise simulation.
Figure 18: How SPSysML would help you in change analysis of a complex CPS?
Figure 19: Clarity of the factor-based rules for integrity improvement.
Figure 20: SPSysML assessment against qualities of great models: 1) Linked to Decision Support, 2) Model Credibility, 3) Clear Scope, 4) V&V With Model, 5) Well-Organized, 6) Analyzable and Traceable, 7) Data Extrapolation, 8) Complete, 9) Internally Consistent, 10) Verifiable, 11) Validation, 12) Model Fidelity, 13) Elegant, 14) Formed for Optimization, 15) Avoid Black Box, 16) Availability of Interfaces, 17) Reusable. The qualities are defined in the questionnaire.
7 Discussion
@@
We state that incorporating SPSysML-based quantitative evaluation of system design indicates significant system structure changes resulting from the proposed quantitative factor-based guidelines applications. An example is FallDetector <<SimAgent>> and FallDetector <<PhyAgent>> integration improvement by a common function isolation as <<SimPhyContSubsys>>. Moreover, the verification shows that 33% of the system’s Physical Agents do not have DTs, and thanks to the design modification, the whole hardware (physical and simulated) is controlled with a common component (DGF=1). This has two main advantages. First, all Digital and Physical Twins pairs share a common control subsystem. Second, if there is a <<PhyAgent>> or <<SimAgent>> without a twin, it can be easily integrated using the embodiment-abstract interface.
@@
We state that the requirement profile for SPSys enables critical analysis and may result in requirement decomposition and explication. An example is the decomposition of the Communication with humans <<SimPhyFunReq>> to TTS and STT <<SimPhyFunReq>> and Dialoge management <<CompFunReq>> during requirement analysis. Thanks to the Dialoge management <<CompFunReq>> separation, the proposed requirement-based procedure for the system composition resulted in the implementation of Talker <<HybAgent>> managing the requirement. As a consequence of the above, the requirement is managed by a component usable in different setups in cooperation with the simulated or physical embodiment. Otherwise, the requirement would be managed separately in simulated and physical embodiments of the robot.
Systems engineering practitioners who reviewed SPSysML confirm the above statements and point out minor difficulties in the SPSysML application.
8 Conclusion
Numerous cyber-physical systems include simulated parts such as Digital Twins (DT), demonstrators, or mockups, which are utilised during their development. We postulate the Simulated-Physical System (SPSys) concept to describe this kind of system. SPSys includes physical, simulated and hybrid parts.
SPSys application is wide; in particular, it can work in a dynamic environment and observe exogenous actions of the inhabitants in the environment. Such a situation is problematic because DT must perfectly mirror its PT. To answer the above needs, we propose a Domain-Specific Language named Simulated-Physical System Modeling Language (SPSysML) that defines SPSys taxonomy and the relationships between the types of SPSys parts.
Integrity and reusability in reliable software development is crucial. Therefore, based on SPSysML terms, we define design evaluation factors that enable quantitative analysis and optimisation targeted to software re-use maximisation between DT and PT in different system setups. We analyze the evaluation factors and show features of the systems that score edge case values of the factors.
Finally, we validate SPSysML, including the evaluation factors and proposed requirement-based system composition in a complex robot system development. We demonstrate step-by-step results of SPSysML application. The validation shows that the requirement profile for SPSys enables critical analysis and may result in requirement decomposition and explication. The validation also demonstrates the positive influence of the SPSysML-based quantitative evaluation on component reusability and integrity between simulated and physical parts.
The presented validation confirms the features distinguishing our method from the others listed in Table 1, e.g.:
  • Quantitative evaluation of integrity between the system’s simulated and physical embodiments,
@@ 8 Conclusion
  • Modeling simulation, physical and hybrid system execution and testing setups,
@@
  • SysML-based specification of systems composed of physical and simulated parts and Digital Twins.
  • Enabling Digital Twins to observe exogenous actions in the simulated world by <<WorldMirrGpAgents>> implementation.
  • Easing requirements tracing using requirement-based system structuring.
@@
  • Requirement-based structuring is a roadmap for SPSys’s design, increasing structural connections between system components and requirements. However, it may not be optimal for any system; therefore, we suggest further iterating the design stage with other system design methods, like those based on the Design Structure Matrix [61].
  • The requirement model defines only types used in the requirement-based structuring method; however, the project team should extend the requirement model if other requirement types are necessary.
  • The integrity evaluation accuracy depends on the system structure granularity, which is the resolution of the evaluation. Therefore, the more detailed the system design is, the more accurate the evaluation.
  • The drawbacks pointed out in the questionnaire. Evaluation factors are objective and abstract from the subjective importance of particular components. However, such aspects can be applied by scaling the proposed factors by importance, priority, etc.
In the future, we plan to automate SPSys structure evaluation with known universal tools like Matlab and Enterprise Architect. Furthermore, this work revealed the need for a CPS structure optimisation indicators taxonomy. SPSysML and future work lead to optimal structure development automation for complex robot systems like dual-arm impedance controlled mobile manipulatiors [10].